OpenStudy (anonymous):

Help? x + 6 times (x - 8) = times (5x - 7) = ?

OpenStudy (anonymous):

So far I've gotten: (x + 6) times (x - 8) = (x + 6)(x - 8) = x^2 - 8x + 6x - 48 = x^2 - 14x - 48 times (5x - 7) = (5x - 7)(x^2 - 14x - 38) = ?

OpenStudy (e.mccormick):

-14x? \((x + 6)(x - 8)\ne x^2 - 14x - 48\)

OpenStudy (anonymous):

Oops, okay. x^2 - 2 - 48?

OpenStudy (e.mccormick):

 Yah.